# Mamba®: Vision Mamba ALSO Needs Registers

### v1

@@ 4 Experiments
Table 3: ImageNet classification results. The throughput is computed on an A100 GPU (numbers with † are taken from (Liu et al., 2024b)). The memory overhead is measured with a batch size of 128 on single GPU. Our results are highlighted in blue.
Model Img. size #Params Throughput Mem. Acc. (%)
@@
ResNet-50 (He et al., 2016) 2242 25M 2388 6.6G 76.2
ResNet-152 (He et al., 2016) 2242 60M 1169 12.5G 78.3
EfficientNet-B3 (Tan & Le, 2019) 3002 12M 496 19.7G 81.6
EfficientNet-B5 (Tan & Le, 2019) 4562 30M 90 78.5G 83.6
EfficientNet-B7 (Tan & Le, 2019) 5602 66M 32 - 84.3
ConvNeXt-T (Liu et al., 2022) 2242 29M 701 8.3G 82.1
ConvNeXt-S (Liu et al., 2022) 2242 50M 444 13.1G 83.1
ConvNeXt-B (Liu et al., 2022) 2242 89M 334 17.9G 83.8
@@
ViT-B/16 (Dosovitskiy et al., 2021) 3842 86M - 63.8G 77.9
ViT-L/16 (Dosovitskiy et al., 2021) 3842 307M - - 76.5
DeiT-S (Touvron et al., 2021) 2242 22M 2390 6.8G 79.8
DeiT-B (Touvron et al., 2021) 2242 86M 1015 14.4G 81.8
DeiT-B (Touvron et al., 2021) 3842 86M 339 63.8G 83.1
@@
VMamba-T (Liu et al., 2024b) 2242 31M 464 7.6G 82.5
VMamba-S (Liu et al., 2024b) 2242 50M 313 27.6G 83.6
VMamba-B (Liu et al., 2024b) 2242 89M 246 37.1G 83.9
@@
Vim-T (Zhu et al., 2024) 2242 7M 1165 4.8G 76.1
Vim-S (Zhu et al., 2024) 2242 26M 612 9.4G 80.5
Mamba®-T 2242 9M 1160 5.1G 77.4
Mamba®-S 2242 28M 608 9.9G 81.1
Mamba®-B 2242 99M 315 20.3G 82.9
3842 94 51.4G 84.3
Mamba®-L 2242 341M 92 55.5G 83.2
3842 28 - 84.5
@@ 4.1 Experimental Settings
[…]
We further evaluate models’ downstream performance on semantic segmentation using the ADE20k (Zhou et al., 2019) dataset, which comprises 150 fine-grained semantic categories distributed across 20K training images, 2K validation images, and 3K test images. Following the existing baseline models (Touvron et al., 2021; Zhu et al., 2024), we choose UperNet (Xiao et al., 2018) as the segmentation head. We utilize AdamW optimizer with a weight decay of 0.01. The models are optimized with a total batch size of 16 for 160k iterations.
@@ 4.2 Main Results
Image classification. As illustrated in Table 3, our Mamba® demonstrates strong performance on ImageNet. Compared to the existing pure Mamba architecture, Vim (Zhu et al., 2024), Mamba® shows a significant improvement, outperforming Vim by 1.3% for the Tiny model and by 0.6% for the Small model. More importantly, compared to Vim, our Mamba® exhibits significant enhancement in scalability: we successfully train a Base (99M parameters, achieving 82.9% accuracy) and even a Large (341M parameters, achieving 83.2% accuracy) size Mamba architectures in vision. This performance can be further enhanced by finetuning with the input resolutions increased to $384\times 384$—our highest accuracy is 84.5%, which outperforms all prior Mamba variants in ImageNet classification.
Semantic segmentation. As shown in Table 5, Mamba® consistently exhibits superior semantic segmentation performance on the ADE20k dataset (Zhou et al., 2019). For example, when compared with Vim-S (Zhu et al., 2024), our Mamba®-S achieves an improvement of 0.4% mIoU. By further scaling up, our Mamba®-B model (featuring 132M parameters) records an mIoU of 47.7%, notably outperforming a similarly-sized DeiT-B model by 2.2% mIoU (results for DeiT are referenced from Liu et al. (2024b)). Additionally, our Mamba®-L (with 377M parameters) also shows great scalability in the segmentation task, achieving 49.1% mIoU on the ADE20k benchmark.
Table 4: Semantic segmentation results on ADE20K. All models are trained with an UperNet head and 512$\times$512 input size. Our results are highlighted in blue
@@
4.3 Ablation Study
Registers’ design choice. Next, we ablate our design choices of registers, i.e., evenly inserting registers and reuse them for final prediction. The results are reported in Table 6. First, we note the performance is sensitive to the positioning of the registers. For instance, positioning all registers at the beginning of the sequence results in a performance decrease of 0.8% (82.1% vs.82.9%). Similarly, positioning all registers in the middle of the sequence, the best strategy reported by Vim (Zhu et al., 2024), still led to a 0.3% drop in performance, which suggests that the sparse distribution of registers helps with feature extraction for Vision Mamba. These noticeable performance gaps highlight the necessity of evenly inserting registers between image tokens, as Mamba’s nature of recurrence makes it sensitive to its token positions in the input sequence.
Further distinctions in register utility are observed when comparing with previous findings (Darcet et al., 2024), which indicate that registers primarily aid in enhancing ViT’s feature representation. In contrast, our study demonstrates that the registers play a crucial role in boosting the quantitative performance of Vision Mamba architectures. Notably, utilizing our default method of evenly distributed registers and reusing all for the final prediction achieved an accuracy of 82.9%, surpassing the approach that uses only one register (R1 only; the rest of the tokens are discarded) by 1.2%. These results affirm that registers constitute a vital component of the Vision Mamba architecture.
Mode Register positions Final prediction Accuracy (%)
Head R1 R2 I1 I2 I3 I4 I5 I6 R1 only 81.3
Mean of registers 81.4
Reduce and concat 82.1
Middle I1 I2 I3 R1 R2 I4 I5 I6 R1 only 81.8
Mean of registers 82.0
Reduce and concat 82.6
Even I1 I2 R1 I3 I4 R2 I5 I6 R1 only 81.7
Mean of registers 82.2
Reduce and concat 82.9
@@ 5 Conclusion
In this paper, we explored the nature of artifacts within the feature maps of Vision Mamba, contrasting these with those observed in Vision Transformers. Specifically, we implemented a novel architecture named Mamba®, incorporating registers strategically to enhance image processing. Our empirical assessments demonstrate that, qualitatively, Mamba® not only reduces the presence of artifacts but also sharpens the focus on semantically relevant areas within images, leading to cleaner and more effective feature maps. Quantitatively, Mamba® not only surpasses its predecessors in accuracy but also exhibits superior scalability, handling larger model sizes with competitive accuracy. We hope this work can establish a solid backbone architecture for future research in optimizing Mamba architectures in vision.
Acknowledge
This work is partially supported by the AWS Cloud Credit for Research program.

### v2

@@ 4 Experiments
Table 3: ImageNet classification results. The throughput is tested on an A100 GPU during training. The memory overhead is measured with a batch size of 128 on a single GPU. Our results are highlighted in blue.
Model Image size Eff. epochs Parameters Throughput Memory Accuracy (%)
@@
ResNet-50 [15] 2242 300 25M 2388 6.6G 76.2
ResNet-152 [15] 2242 300 60M 1169 12.5G 78.3
EfficientNet-B3 [41] 3002 300 12M 546 19.7G 81.6
EfficientNet-B5 [41] 4562 300 30M 143 78.5G 83.6
EfficientNet-B7 [41] 5602 300 66M 61 $>$80G 84.3
ConvNeXt-T [29] 2242 300 29M 635 8.3G 82.1
ConvNeXt-S [29] 2242 300 50M 412 13.1G 83.1
ConvNeXt-B [29] 2242 300 89M 305 17.9G 83.8
@@
ViT-B/16 [8] 3842 300 86M 201 63.8G 77.9
ViT-L/16 [8] 3842 300 307M 95 $>$80G 76.5
DeiT-S [42] 2242 300 22M 1924 6.8G 79.8
DeiT-B [42] 2242 300 86M 861 14.4G 81.8
DeiT-B [42] 3842 300 86M 201 63.8G 83.1
DeiT-III-B [43] 2242 315 86M 861 14.4G 83.2
@@
VMamba-T [27] 2242 300 31M 464 7.6G 82.5
VMamba-S [27] 2242 300 50M 313 27.6G 83.6
VMamba-B [27] 2242 300 89M 246 37.1G 83.9
LocalVMamba-S [19] 2242 300 50M 289 28.0G 83.7
PlainMamba-L3 [50] 2242 300 50M 256 28.5G 82.3
@@
Vim-T [54] 2242 300 7M 750 4.8G 76.1
Vim-S [54] 2242 300 26M 395 9.4G 80.5
LocalVim-S [19] 2242 300 28M 404 9.6G 81.2
Mamba®-T 2242 300 9M 746 5.1G 77.4
Mamba®-S 2242 230 28M 391 9.9G 81.4
Mamba®-B 2242 230 99M 196 20.3G 83.0
3842 230 99M 63 51.4G 84.3
Mamba®-L 2242 230 341M 67 55.5G 83.6
3842 230 342M 23 $>$80G 84.5
@@ 4.1 Experimental Settings
We primarily evaluate our Mamba® on the standard ImageNet [7] dataset, which consists of $\sim$1.28 million training images and 50,000 validation images spread across 1,000 categories. Our training setup mostly follows the protocols established in DeiT [42]. Specifically, we use AdamW optimizer [30] with a momentum of 0.9, a weight decay of 0.05, a cosine annealing learning rate starting at 1$\times$10-3, a batch size of 1024 for Mamba®-Tiny, Small, and Base, and a batch size of 2048 for Mamba®-Large. For better efficiency and preventing over-fitting, we train each model for 300 epochs in 128$\times$128 input size and fine-tune in 224$\times$224 with stronger data augmentation strategies. We also empirically find a 100-epoch intermediate finetuning with weak augmentation can further improve the results. In total, the training process leads to $\sim$230 effective training epochs in 224$\times$224 image size, yet significantly outperforms its 300-epoch counterparts. A detailed training recipe can be found in the Appendix.
We further evaluate models’ downstream performance on semantic segmentation using the ADE20k [52] dataset, which comprises 150 fine-grained semantic categories distributed across 20K training images, 2K validation images, and 3K test images. Following the existing baseline models [42, 54], we choose UperNet [49] as the segmentation head. We utilize AdamW optimizer with a weight decay of 0.01. The models are optimized with a total batch size of 16 for 160k iterations.
@@ 4.2 Main Results
Image classification. As illustrated in Table 3, our Mamba® demonstrates strong performance on the ImageNet-1k classification benchmark. Compared to the existing pure Mamba architecture, Vim [54], Mamba® shows a significant improvement, outperforming Vim by 1.3% for the Tiny model and by 0.6% for the Small model. More importantly, compared to Vim, our Mamba® exhibits significant enhancement in scalability: we successfully train a Base (99M parameters, achieving 83.0% accuracy) and even a Large (341M parameters, achieving 83.2% accuracy) size Mamba architectures in vision, demonstrating that Mamba-based causal models can also learn with hundreds of million parameters in image modeling. This performance can be further enhanced by finetuning with the input resolutions increased to $384\times 384$—our highest accuracy is 84.5%, which outperforms all prior Mamba variants in ImageNet classification.
Semantic segmentation. As shown in Table 5, Mamba® consistently exhibits superior semantic segmentation performance on the ADE20k dataset [52]. For example, when compared with Vim-S [54], our Mamba®-S achieves an improvement of 0.4% mIoU. By further scaling up, our Mamba®-B model (featuring 132M parameters) records an mIoU of 47.7%, notably outperforming a similarly-sized DeiT-B model by 2.2% mIoU (results for DeiT are referenced from [27]). Additionally, our Mamba®-L (with 377M parameters) also shows great scalability in the segmentation task, achieving 49.1% mIoU on the ADE20k benchmark.
4.3 Ablation Study
Mode Register positions Final prediction Accuracy (%)
Head R1 R2 I1 I2 I3 I4 I5 I6 R1 only 81.3
Mean of registers 81.4
Reduce and concat 82.1
Middle I1 I2 I3 R1 R2 I4 I5 I6 R1 only 81.8
Mean of registers 82.0
Reduce and concat 82.6
Even I1 I2 R1 I3 I4 R2 I5 I6 R1 only 81.7
Mean of registers 82.2
Reduce and concat 83.0
Table 5: Semantic segmentation results on ADE20K. All models are trained with an UperNet head and 512$\times$512 input size.
@@ 4.3 Ablation Study
Registers design choice. Next, we ablate our design choices of registers, i.e., evenly inserting registers and reuse them for final prediction. The results are reported in Table 4. First, we note the performance is sensitive to the positioning of the registers. For instance, positioning all registers at the beginning of the sequence results in a performance decrease of 0.8% (82.1% vs. 83.0%). Similarly, positioning all registers in the middle of the sequence, the best strategy reported by Vim [54], still led to a 0.3% drop in performance, which suggests that the sparse distribution of registers helps with feature extraction for Vision Mamba. These noticeable performance gaps highlight the necessity of evenly inserting registers between image tokens, as the recurrent formulation makes Mamba sensitive to positions in the sequence.
@@
Further distinctions in register utility are observed when comparing with previous findings [6], which indicate that registers primarily aid in enhancing ViT’s feature representation. In contrast, our study demonstrates that the registers play a crucial role in boosting the quantitative performance of Vision Mamba architectures. Notably, utilizing our default method of evenly distributed registers and reusing all for the final prediction achieved an accuracy of 83.0%, surpassing the approach that uses only one register (R1 only; the rest of the tokens are discarded) by 1.2%. These results affirm that registers constitute a vital component of the Vision Mamba architecture.
@@ 5 Conclusion
In this paper, we explored the nature of artifacts within the feature maps of Vision Mamba, contrasting these with those observed in Vision Transformers. Specifically, we implemented a novel architecture named Mamba®, incorporating registers strategically to enhance image processing. Our empirical assessments demonstrate that, qualitatively, Mamba®not only reduces the presence of artifacts but also sharpens the focus on semantically relevant areas within images, leading to cleaner and more effective feature maps. Quantitatively, Mamba®not only surpasses its predecessors in accuracy but also exhibits superior scalability, handling larger model sizes with competitive accuracy. We hope this work can establish a solid backbone for future research in optimizing Mamba architectures in vision.
Acknowledgment
This work is supported by NEI Award ID R01EY037193 and ONR N00014-23-1-2641. We also appreciate the computation resources provided by the AWS Cloud Credit for Research program.
Appendix A: Technical Details
Table 7: Pre-training configurations
Configuration Small Base Large
input size 128
epochs 300
optimizer AdamW
weight decay 0.05
base learning rate 5e-4 2e-4 2e-4
batch size 1024 2048 2048
drop path 0.1
label smoothing
random erasing
Rand Augmentation
repeated augmentation
ThreeAugmentation
Table 8: Intermediate training configurations
Configuration Small Base Large
input size 224
epochs 100
optimizer AdamW
weight decay 0.05
base learning rate 2e-4
batch size 1024
drop path 0.2 0.4 0.4
label smoothing
random erasing
Rand Augmentation
repeated augmentation
ThreeAugmentation
Table 9: Fine-tuning configurations
Configuration Small Base Large
input size 224
epochs 20
optimizer AdamW
weight decay 0.1
base learning rate 1e-5
batch size 512
drop path 0.2 0.4 0.6
label smoothing 0.1
random erasing
Rand Augmentation rand-m9-mstd0.5-inc1
repeated augmentation
ThreeAugmentation
We train Mamba®-Tiny by the configurations of DeiT-Tiny [42] but follow a weaker data augmentation strategy used in [43]. For bigger sizes of Mamba®models, we use a three-stage training approach to prevent over-fitting and reduce effective training epochs. We summarize the recipes of pre-training, intermediate training, and fine-tuning in Table 7, Table 8, and Table 9, respectively. For all stages, the learning rate is calculated by base lr $\times$ batch size / 512, following a cosine decay scheduling with 5 epochs warmup. We use color jitter with a factor of 0.3, mixup and cutmix with alpha setting to 0.8 and 1.0, respectively.
Appendix B: Theoretical Justification
We demonstrate that Mamba® performs well because it effectively addresses the vanishing gradient problem which broadly appears in recurrent models. This problem is similar to the gradient propagation issue in training deep networks without residual connections, but occurs along the sequence dimension. Here we derive it with a simplified formulation ${\bm{y}}_{t}={\bm{F}}({\bm{y}}_{t-1};\theta)$ where ${\bm{y}}$ is output and $\theta$ denotes the parameters of the SSM ${\bm{F}}$. We have
$d{\bm{y}}_{t}$ $=\nabla_{\theta}{\bm{F}}({\bm{y}}_{t-1})d\theta+\nabla_{\bm{y}}{\bm{F}}({\bm{y}}_{t-1})d{\bm{y}}_{t-1}$
$=\nabla_{\theta}{\bm{F}}({\bm{y}}_{t-1})d\theta+\nabla_{\bm{y}}{\bm{F}}({\bm{y}}_{t-1})(\nabla_{\theta}{\bm{F}}({\bm{y}}_{t-2})d\theta$
$+\nabla_{\bm{y}}{\bm{F}}({\bm{y}}_{t-2})d{\bm{y}}_{t-2})=\cdots$
$=(\nabla_{\theta}{\bm{F}}({\bm{y}}_{t-1})+\nabla_{\bm{y}}{\bm{F}}({\bm{y}}_{t-1})\nabla_{\theta}{\bm{F}}({\bm{y}}_{t-2})+\cdots)d\theta.$
Given a loss $L({\bm{y}}_{T})$ applied to the $T$-th token, we have $dL=\nabla_{\bm{y}}L({\bm{y}}_{T})d{\bm{y}}_{t}$ and gradient of parameters $\Delta\theta$ equals
$\nabla{\bm{y}}L({\bm{y}}_{T})(\nabla_{\theta}{\bm{F}}({\bm{y}}_{t-1})+\nabla_{\bm{y}}{\bm{F}}({\bm{y}}_{t-1})\nabla_{\theta}{\bm{F}}({\bm{y}}_{t-2})+\cdots),$
where repeated multiplications of gradients appear:
$\nabla_{\bm{y}}{\bm{F}}({\bm{y}}_{t-1})\nabla_{\bm{y}}{\bm{F}}({\bm{y}}_{t-2})\nabla_{\bm{y}}{\bm{F}}({\bm{y}}_{t-3})\cdots$
As derived, the number of repeated gradient multiplications depends on the number $T$. Here we show a simple case of Mamba® with nine image tokens and two registers. As we employ a reduce-and-concat module, the two registers are bound together by ${\bm{y}}_{r2}=g({\bm{y}}_{r1})$. The gradient propagation from ${\bm{y}}_{r2}$ to ${\bm{y}}_{1}$ can be written as
$d{\bm{y}}_{r2}=\nabla_{{\bm{y}}_{1}}{\bm{F}}({\bm{y}}_{r2})d{\bm{y}}_{1}+\nabla_{{\bm{y}}_{r1}}{\bm{F}}({\bm{y}}_{r2})\nabla_{{\bm{y}}_{1}}{\bm{F}}({\bm{y}}_{r1})d{\bm{y}}_{1},$
where the second term only involves four gradient multiplications. If ${\bm{y}}_{r1}$ and ${\bm{y}}_{r2}$ are not bound together by $g$, the number of gradient multiplications becomes eight.